a club member has 12 members. in how many ways can a president, vice president, and a secretary be chosen?... is it 12X11X10
Answer by edjones(8007)

Permutation of n things taken r at a time: nPr=n!/(n-r)!
12P3=12*11*10*9!/9!=12*11*10=1320 ways.
.
Ed